OpenStudy (anonymous):

Hey guys I will give a medal and like if someone checks my work For the line segment whose endpoints are L (0, 1) and M (2, 8), find the y value for the point located 3 over 5 the distance from L to M. 5.2<--------------------my answer 3.5 4.8 1.6

OpenStudy (anonymous):

....

OpenStudy (anonymous):

@triangleguru

OpenStudy (anonymous):

Yes; answer is OK: y = yL + 3/5 dy = 1+ 3/5 * (8-1) = 5.2

OpenStudy (anonymous):

Thanks, di you mind helping me with others if thats ok? @triangleguru